OpenStudy (anonymous):

A pole that is 25 feet tall casts a shadow 65 feet long. At the same time, another pole casts a shadow having a length of 39 feet. What is the height, in feet, of the second pole? Answer 10.5 14 15 41.7

jimthompson5910 (jim_thompson5910):

(Height of pole1)/(Shadow length of pole1) = (Height of pole2)/(Shadow length of pole2) 25/65 = x/39 25*39 = 65x 975 = 65x 975/65 = x 15 = x x = 15 So the length of the second pole is 15 feet.